Today has been a busy day for the Crew 185. As it was our last day before entering simulation, we needed to prepare the hab in the best possible way and to fix as many things as possible before closing our doors.
We decided to work in two groups. Ilaria and David took the car to buy a new battery for the generator, soil for the greenhab and food. Arno, Thibault and John set the laboratory with the 3D printer, the medical supplies, the X-1 spacesuit prototype, etc…
This afternoon, we repaired the generator by replacing the battery and changing a part of the cooling liquid and the oil. It is working better at the moment when I write these lines! Hopefully, it will keep being in a good shape.
We decided that we would begin the simulation tomorrow morning when waking up. Mars, here we come!
The Crew 185
Written by David Murray
Supervised by Trisha Randazzo
December 16, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient (no heating or/cooling)
Ambient with window/door open
· Outside temperature during work hours: -4.5 C
· At 20:25 inside temperature was 15 C.
· Inside temperature high 31
· Inside temperature low 15
· Inside humidity: 16%
· Functioning nominally
· Turns on at 16 C stops at 20 C
· Not available
Both heating and cooling
Shade cloth on/off
Average temperature: (N/A until new sensor is delivered)
Changes to crops: Note all emergence of seedlings, death of seedlings, etc
· Two of the bean plants are still drooping
· The sprouting bean plants now has 4 flowers
Daily water usage for crops:
· Used more water to accommodate the drier planter box.
Time(s) of watering for crops:
Morning research observations: (Currently not operational)
Changes to research plants:
Daily watering and amount of water used:
· 4 Gallons
Aquaponics: (Currently not operational)
Narrative: Any other information you want to share
· Completed crew handover activities for GreenHab. David is now well trained as the new GreenHab Officer.
· Soil and a fire extinguisher
Name: Thomas Horn Crew: 184
Sky Conditions: Clear
Wind Conditions: Light
Observation Start Time: 2:00 pm
Observation End Time: 2:30 pm
Summary: No observing, shutdown Observatory and returned Laptop to HAB
Objects Viewed: None
Problems Encountered: None
MDRS Sol Summary Report for Sol 14
Summary Title: Signing off!
It’s been a please working with all of you, and thanks for your help.
It was a fitting send off to our team that our day started with a
generator failure. Upon our morning rounds after waking up it was
noticed the Generator had failed with an ‘Underspeed’ Fault. We were
happy to see that SOC was holding steady at 100% (thank you solar
arrays!) but this was a worrying sign, especially as we didn’t want to
hand over a crippled ship to the oncoming crew.
Investigation into the generator showed that the loose connections
that were repaired previously were still holding strong, so there was
no immediate cause of the generator fault. After attempting to
restart the generator a low battery voltage was observed (reading
7.6volts, supposed to be 9.6 volts). It appears that after the
generator failed it drained the battery down displaying the fault
message on its status plate. After a jumpstart from our friendly
neighbors we were able to get the generator back up and running. This
issue is likely to re-occur as we are still not sure of the original
cause. Our next step will be to Remove and Replace the generator
batteries, fingers crossed this fixes the problem. Best of luck to
future crews in continuing to troubleshoot this issue.
Besides that morning excitement the day was one of cleaning and
packing, with a short break in the afternoon to do a short EVA around
the HAB for some last minute filming to get the shots we’ve missed so
far this mission. We hope to leave this place better than we found
This report comes to you with the assistance of Mission Support as our
Internet allowance has run out. Sometime today the same laptop that
killed our bandwidth before had WiFi turned On, after which it sucked
our bandwidth dry again. Still don’t know why this particular laptop
uses bandwidth at such a prodigious rate, but we’ve turned it off
again and comms should be returned to normal.
Hope to work with you all again in the future.
Ad Astra Per Ardua,
Sol Activity Summary:
1. Filming EVA
2. Cleaning / Packing
4. Reports and general HAB maintenance / checks
6. Generator Failure and Jumpstart
Look Ahead Plan:
1. Crew Handover Training
Anomalies in work:
1. HAB Leak from under EVA door continuing (per Mission Control
leak is expected and has been present for months, presumed to be from
2. Generator Underspeed Fault
Weather: Cold, but temps are above freezing in mid-day
Crew Physical Status: All crew in good physical health.
EVA: None today.
Reports to be filed: